OpenStudy (anonymous):

An ac circuit consists of a voltage source v=200sin120πt and a 6μF capacitor in series. Calculate the current established in the circuit

OpenStudy (shamim):

can i give u an attachment for ur solution

OpenStudy (shamim):

u r most welcome

OpenStudy (anonymous):

That's another, kinda similar...

OpenStudy (anonymous):

You should be able to solve this one on your own then :P

OpenStudy (anonymous):

If any voltage source has frequency f, and amplitude Vo what is the expression for instantaneous velocity?!

OpenStudy (anonymous):

i mean instantaneous voltage?

OpenStudy (anonymous):

\[V=V _{0}Sin(2 \pi ft)\]

OpenStudy (anonymous):

so you have the frequency.. the only thing you need to evaluate is Vo.. the voltmeter reads 240V.. which has to be an AC voltmeter THOSE ARE YOU CLUES :P

OpenStudy (shamim):

i think Vrms=240volt.

OpenStudy (anonymous):

@shamim congratulations.. you gave away the most important part of the question .. !! :-/ !

OpenStudy (shamim):

Vrms=0.707Vo

OpenStudy (anonymous):

@shamim Thats enough :O.. let him do the problem :O :O : O!!

OpenStudy (shamim):

ok @Mashy

OpenStudy (shamim):

thanks @d3Xter

OpenStudy (anonymous):

you are kidding right o.O

OpenStudy (anonymous):

\[I = \frac{ d }{ dt}\left\{ 6 \times 10^{-6} \times 200\sin (120 \pi t) \right\}\]

OpenStudy (anonymous):

i don't understand.. what exactly do you have to find out?

OpenStudy (anonymous):

and how did you get that expression?

OpenStudy (anonymous):

From \[Q=CV\]

OpenStudy (anonymous):

ok ok.. just take the derivative!

OpenStudy (anonymous):

Then \[I = \frac{ dQ }{ dt }\]

OpenStudy (anonymous):

I've done that already.

OpenStudy (anonymous):

ok.. and that is not the correct answer?

OpenStudy (anonymous):

How do I simplify further

OpenStudy (anonymous):

there is no further simplification.. that is the expression for instantaneous current o.O

OpenStudy (anonymous):

\[I=0.144 \pi \cos120 \pi t\]

OpenStudy (anonymous):

you can't further simplify that :O

OpenStudy (anonymous):

Oh. that is the RMS value .. the DC equivalent!.. stupid teacher should have mentioned that in the question.. go all gangstaaa on your teacher xD